A gas in a closed container is added with 20J of heat. The lid to the container is then lifted for 20 cm with 2.5 N of force. Calculate how much the energy added to the system.


1
Expert's answer
2022-01-18T13:42:47-0500

"Q=U+A=U+Fl=20+2.5\cdot 0.2=20.5~J."